angle A and angle B are complimentary. if angle a is 6 more that 3 times angle b what is each angle. solve in 5 steps
Answer by katealdridge(100)

A+B=90. A=3B + 6
Substitute 3B+6 for A in the first equation.
3B+6+B=90 combine like terms
4B+6=90 subtract 6 from both sides
4B=84 divide both sides by 4
B=21
A=90-21=69.